# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v12

@@ 2.3 Mining Technology and Pollutions
Pollution may be another incidence that could influence the regional economy. Generally, industrial pollution was not recognized as a common social problem in pre-war Japan. As a result, there is little documentation of pollution due to coal mining. [^16] An exception is the Ashio Copper Mine Incident in the late 19th century (e.g., Notehelfer 1975). In the field of development economics, von der Goltz and Barnwal (2019) provides evidence that mining is associated with child stunting among today’s developing nations. Anecdotes, however, suggest that there are a few potential pathways that generate pollution around coal mines.
First was the soot and smoke caused by coal consumption to run steam-winding machines. Steam-winding engines were introduced early in the mechanization of the hauling process and exhausted the smoke from burning coal in the boiler. However, these steam-powered winders were gradually replaced with electric winders during the interwar period (Japan Mining Association 1930, 1931). Thus, this channel might be unlikely to impact the regional population health. Second, presumably more relevant, is pollution from wastewater generated during mining. As coal mining became increasingly mechanized, some mines began introducing water-washing machines during coal sorting. Unfortunately, there is little historical documentation systematically describing pollution caused by wastewater from coal mines. However, coal sludge from coal selection had contaminated rivers to some extent at that time. An example is Onga river in the Chikuhō coalfield, which was called a “zenzai” (sweet red-bean soup) because the water was contaminated by the liquid waste (Chiba and Yamada 1964). Coal sludge contains several metal toxicities, including mercury (Hg) and cadmium (Cd), which are known to be associated with the incidence of miscarriages and stillbirths (Amadi et al. 2017; Sergeant et al. 2022). This suggests the possibility that early-life mortality would have been higher in the coal mining region with greater river accessibility. From the 1920s, however, some mines installed settling ponds to purify wastewater and circulate it to coal cleaning (Advisory Council on Mining 1932, pp. 313–316). This might work to mitigate the pollution risks around the mines due to the wastewater.
@@ 3 Conceptual Framework
The fertility decision in response to the regulation of female miners described in the previous section would affect the population in the mining area. The decline in the relative wage rate of females due to the regulation decreases the opportunity cost of housework, wheras it also decreases their earnings in the labor markets. While the former mechanism may lead to higher fertility through the substitution effect from consumption, the latter may result in lower fertility via the income effect. This section offers a brief overview of a theoretical framework predicting the relevant mechanism given the historical evidence about the coal mines in prewar Japan. To do so, I consider the collective model proposed by Siegel (2017) that illustrates the optimal fertility based on the couple household’s utility maximization. [^17] For the sake of brevity, I consider a simplified model by abstaining the utility from leisure and home production parameters. This offers a materially similar interpretation of the optimal fertility choice to the original generalized model. See Browning et al. (2014) for a comprehensive review on the cooperative and non-cooperative models of the couples.
Let us consider that both males and females derive utilities from consumption ($c$) and having children ($b$) under additively separable utility function as $u_{g}(c_{g},b)=\log c_{g}+\tau\log b$, where $g\in\{m,f\}$ indicates a spouse. Raising children requires a home production $y(b)=\pi b$ with $\pi>0$, which is accomplished using labor-saving technologies available in the markets ($s$) and parents’ total home labor input ($\Omega$) as: $y(b)=s^{\gamma}\Omega^{1-\gamma}$. $\gamma$ represents the availability of childcare services in the economy ($10\leq\gamma<$ 1). The parents’ total home labor input is specified in the constant elasticity of the substitution function as: $\Omega=\left(\kappa_{m}h_{m}^{1-\rho}+\kappa_{f}h_{f}^{1-\rho}\right)^{\frac{1}{1-\rho}}$, where $\kappa_{g}$ is the productivity of housework including childcare, $h_{g}\geq 0$ is the home labor input, and $\rho>0$ is the inverse of the elasticity of substitution. The labor income is derived as the product between wage $w_{g}$ and labor input $n_{g}\geq 0$, which is used for their consumption and for purchasing home labor-saving services.
The representative household maximizes the sum of utilities $\theta u_{m}(c_{m},b)+(1-\theta)u_{f}(c_{f},b)$ subject to the budget constraints $c_{m}+c_{f}+s=w_{m}n_{m}+w_{f}n_{f}$ and time constraint $h_{g}+h_{g}=1~{}\text{for}~{}g\in\{m,f\}$. The optimal fertility choice ($b^{*}$) is then given in a closed-form function. One can confirm that the sign of the derivative of the optimal fertility with respect to the wage of female ($\frac{\partial b^{*}}{\partial w_{f}}$) is negative under the following condition:
@@
Online Appendix B provides details of the derivation. Under a paternalistic social norm stipulated in the Old Civil Code, the head’s housework productivity, particularly childrearing, shall be negligible ($\kappa_{m}=0$). It follows that condition 1 implies $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$ for my empirical setting. Since the marketization of childcare ($\gamma$) was considerably low at that time, the optimal fertility is negatively associated with the relative wage ($w_{f}/w_{m}$). [^18] As mechanization progressed, miners were required to accumulate knowledge and experience in operating machinery. Management, therefore, had an incentive to encourage miners to form families and continue employment over a long term (Ogino 1993). For example, many coal mines were equipped with nurseries and other welfare facilities from the 1920s (Ministry of Commerce and Industry 1926). However, evidence shows that these nurseries required fees and had covered only a small part of their childcares. See Online Appendix B for details. This result suggests the following two predictions:
@@
  1. Prediction 1: Before the female labor regulation, the mining area could have lower fertility, given the higher relative wages in the mining sector.
@@
  2. Prediction 2: After the female labor regulation, the mining area could have recovered fertility in response to the reduction in the relative wages.
@@
The short-run population dynamics in a local economy depend on natural and social changes. [^19] From year $t$ to $t+1$, this geometric population growth model is specified as $\Psi_{t+1}-\Psi_{t}=(\Upsilon_{t}-\Omega_{t})+\Pi_{t}$, where $\Psi$, $\Upsilon$, $\Omega$, and $\Pi$ indicate the number of people, live births, deaths, and net migrants, respectively. In the former change, the lower fertility under the substitution effect before the regulation would not be a major factor in population growth in the mining area. In turn, the fertility gain after the regulation may push up the growth. Mortality term in the natural change is given outside of the household’s fertility choice in the short-run. The regulation of female miners shall reduce their occupational hazards, which may improve female mortality. The improvements in maternal health may also mitigate prenatal and neonatal deaths (Section 2.3). Both ameliorations would work to reduce population losses in the mining area. The influence of social change on local population growth is uniquely determined when the sign of natural change is determined. Moreover, social change shall be positive in the mining area because the regional wage gap would encourage the influx of workers from surrounding agrarian areas into the mining area. For female workers, the wage reduction via the regulation may reduce the attractiveness of coal mining labor. However, it may not offset the benefits for couples moving into the mining area because male miners’ wages were not cut off by the regulation. [^20] In other words, since female workers migrated with their husbands, the migration of couples is not affected by the regulation as much as that of single workers. To summarize, while the mining area had an influx of workers throughout, the fertility and mortality in the mining area could be affected by the enforcement of the revised regulation of 1933.
@@ 4 Data
This study created a unique dataset of the demographics across mining areas in the Japanese archipelago. Because mining is a localized economic activity, a smaller lattice dataset is preferred to identify the impact of mines. I collected and digitized a set of official census-based municipal-level statistics published in the 1930s.
@@ 4.1 Mine Deposit
I use official reports named Zenkoku kōjyō kōzan meibo (lists of factories and mines) published by Association for Harmonious Cooperation (1932, 1937) (hereafter, called the AHC), which document coal mine locations measured in October 1931 and October 1936. The AHC listed all mines with $50$ miners and more. Thus, the main target of this study is the average impact of small- to large-scale coal mines and not the very small-scale collieries with fewer than $50$ mimers. [^21] Generally, a few coal mines are clustered in a coal mining municipality. For example, 85 (108) out of 93 (115) municipalities with mines had one to four mines in the 1930 (1935) sample. In this sense, this study estimates the average impacts of these general coal mining municipalities. In Online Appendix D.11, I confirmed that my main findings are not sensitive to a small number of municipalities with several mines. Despite this, the AHC is still comprehensive as it includes information on small-scale mines, which have been neglected in the literature.
I define the treatment group as the municipalities located within 0–5 km of the centroid of a municipality with mines. [^22] I used an official shapefile provided by the Ministry of Land, Infrastructure, Transport and Tourism for geocoding. See Online Appendix C.1 for the details. A 5 km threshold is fundamentally plausible because the mean value of the distance to the nearest neighborhood municipality is approximately 4 km. In addition, the estimated effects disappeared in the statistical sense outside the 5 km range, which means that the potential spillover effect is captured under this threshold. Online Appendix D.6 summarizes finer details of the validity of the threshold. For the control group, I set the threshold as 5–30 km from the centroid. [^23] My definition of both groups is conservative compared with that in the literature, which uses 10–20 km and 100–200 km as the thresholds of the treatment and control groups, respectively (Wilson 2012; Aragón and Rud 2015; Kotsadam and Tolonen 2016; Benshaul-Tolonen 2019). This is because Japan has a smaller and thinner archipelago than countries studied in the literature, such as South Africa (Online Appendix Figure 2). In Online Appendix D.6, I further provide evidence that 5–30 km is a plausible distance for the control group given that the impacts of the coal mines are concentrated within 5 km distance from the mines.
@@
Figure 3: Location of Coal Mines around Chikuhō Coalfield in Kyushū
Notes: 1. The white circles indicate the location of municipalities with coal mines around Chikuhō coalfield (Figure 2a) in Kyushū region. Figure 3a shows the coal mines in Chikuhō, Kasuya, and Miike coalfields in Fukuoka prefecture, Karatsu coalfield in Saga prefecture, and Miike coalfield in Kumamoto prefecture. Amakusa coalfield in Kumamoto prefecture is added in Figure 3b (on the lower left of the figure). Ōita prefecture does not have coal mines but is shown in the figures to explain the border of the sample.
2. Treatment (control) group includes municipalities within 5 (between 5 and 30) km from a mine. The excluded municipalities are shown as empty lattices in the figures.
@@
Figure 2 indicates the coal mine locations from the AHC and the corresponding treatment and control groups by year. Most coal mines are located in three specific areas in the southernmost, central, and northeastern regions: these are representative coalfields called Chikuhō, Jyōban, and Ishikari-Kushiro Coalfields, respectively. Notably, a comparison of Figures 2a and 2b shows that these agglomerations were stable over time. For a closer look, Figure 3 provides an example map for the most representative coalfield called Chikuhō Coalfield in Kyushū, Japan’s southernmost island (indicated in Figure 2a). One can see that the distribution of mines has changed little; only a few mines were opened or closed between 1931 and 1936. This feature of coal mines makes it difficult to use within-variation in the spatiotemporal distribution of the mines for identification. Online Appendix D.8 discusses this point in detail. [^24] In short, it is infeasible to implement a difference-in-differences estimator for my full sample because of the lack of within-variation (i.e., the number of municipalities that experienced the opening/closing of the coal mines). Nevertheless, in Online Appendix D.8, I will show the results from the two-group by two-period difference-in-differences setting for a small subsample including Ogawa and Miyoshiyama coal mines (Figure 2b), which provide the results reasonable to my findings.
Finally, I explain the generation of the analytical samples. From $11,151$ total municipalities, I first retain the municipalities within a $30$ km radius from the centroid of the municipalities with mines. I then exclude the intersection of the different types of mines, such as municipalities within $5$ km from a gold, silver, and copper (GSC) mine. Online Appendix Table C.3 presents the summary statistics of the treatment variables. The number of municipalities in the analytical sample based on the 1931 and 1936 AHC were $1,140$ and $1,364$, respectively. Meanwhile, the proportion of treated municipalities was approximately $15$% and $13$%, respectively, for each sample. [^25] The number of municipalities with coal mines in each analytical sample is 93 and 115, respectively. The slight reduction in the share of the treatment group is due to the increase in the number of enclaves, which I explained above. A mine opening in the enclave area increases the number of controlled (i.e., surrounding) municipalities, which leads to a decrease in the relative share of the treatment group. See, for example, Ogawa and Miyoshiyama coal mines in Tōhoku region (Figure 2b).
@@ 4.2 Geological Stratum
The spatial distribution of a specific geological stratum created in the Cenozoic era, which includes some carboniferous ages, is used as an instrumental variable (IV) for the location of coal mines (Section 5.1). Data on the geological strata are obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism, which includes roughly nine thousand stratum points. I matched each municipality in my analytical sample with the nearest stratum point to identify the municipalities with the relevant stratum. Online Appendix C.2 provides finer details of the definition and an example of a geological columnar section of a representative coal mine to show the relevance of the IV. Online Appendix Table C.3 presents the summary statistics of the indicator variables for the stratum.

### v14

@@ 2.3 Mining Technology and Pollutions
Pollution may be another factor that could influence the regional economy. Generally, industrial pollution was not recognized as a common social problem in pre-war Japan. Consequently, there is little documentation of pollution due to coal mining. [^17] An exception is the Ashio Copper Mine Incident in the late 19th century (e.g., Notehelfer 1975). In the field of development economics, von der Goltz and Barnwal (2019) provides evidence that mining is associated with child stunting in nations that are still economically developing. However, anecdotes suggest that a few potential pathways generate pollution in coal mines.
First was the soot and smoke caused by coal burning to run steam-powered winding machines. Steam-powered winding engines were introduced early in the mechanization of the hauling process (Section 2.1) and exhausted smoke generated from burning coal in the boiler. However, these steam-powered winders were gradually replaced by electric winders during the interwar period (Japan Mining Association 1930, 1931). Thus, this channel may be unlikely to impact the regional population health. Second, presumably more relevant, is pollution from wastewater generated during mining. As coal mining became increasingly mechanized, some mines began introducing water-washing machines during coal sorting. Unfortunately, there is little historical documentation systematically describing pollution caused by wastewater from coal mines. However, coal sludge from coal selection started contaminating rivers at the time. An example is Onga river in the Chikuhō coalfield, which was called a “zenzai” (sweet red-bean soup) because the water was contaminated by the liquid waste (Chiba and Yamada 1964). Coal sludge contains several metal toxins, including mercury (Hg) and cadmium (Cd). These are known to be associated with the incidence of miscarriages and stillbirths (Amadi et al. 2017; Sergeant et al. 2022). This suggests the possibility that early-life mortality would have been higher in the coal mining regions with greater river accessibility. From the 1920s, however, some mines installed settling ponds to purify wastewater and recirculate it during coal cleaning (Advisory Council on Mining 1932, pp. 313–316). This could mitigate the pollution risks around the mines due to the wastewater.
@@ 3 Conceptual Framework
The regulation of female miners’ working conditions described in the previous section would influence the fertility in the mining area. The decline in females’ relative wage rate due to regulation can decrease the opportunity cost of housework, whereas it can also decrease their earnings in the labor market. While the former mechanism could lead to higher fertility through the substitution effect from consumption, the latter may result in lower fertility via the income effect. This section offers a brief overview of a theoretical framework predicting the relevant mechanism given the historical evidence about the coal mines in prewar Japan. To do so, I consider the collective model proposed by Siegel (2017) that illustrates the optimal fertility based on the couple household’s utility maximization. [^18] For the sake of brevity, a simplified model that abstains the utility from leisure and home production parameters is considered. This offers a materially similar interpretation of the optimal fertility choice to the original generalized model. See Browning et al. (2014) for a comprehensive review on the cooperative and non-cooperative models of the couples.
Let us consider that both males and females derive utilities from consumption ($c$) and having children ($b$) under an additively separable utility function as $u_{g}(c_{g},b)=\log c_{g}+\tau\log b$, where $g\in\{m,f\}$ indicates a spouse. Raising children requires a home production $y(b)=\pi b$ with $\pi>0$, which is accomplished using labor-saving technologies available in the markets ($s$) and parents’ total home labor input ($\Omega$) as: $y(b)=s^{\gamma}\Omega^{1-\gamma}$. $\gamma$ represents the availability of childcare services in the economy ($10\leq\gamma<$ 1). The parents’ total home labor input is specified in the constant elasticity of the substitution function as: $\Omega=\left(\kappa_{m}h_{m}^{1-\rho}+\kappa_{f}h_{f}^{1-\rho}\right)^{\frac{1}{1-\rho}}$, where $\kappa_{g}$ is the productivity of housework including childcare, $h_{g}\geq 0$ is the home labor input, and $\rho>0$ is the inverse of the elasticity of substitution. The labor income is derived as the product between wage $w_{g}$ and labor input $n_{g}\geq 0$, which is used for their consumption and for purchasing home labor-saving services.
The representative household maximizes the sum of utilities $\theta u_{m}(c_{m},b)+(1-\theta)u_{f}(c_{f},b)$ subject to budget constraints $c_{m}+c_{f}+s=w_{m}n_{m}+w_{f}n_{f}$ and time constraint $n_{g}+h_{g}=1~{}\text{for}~{}g\in\{m,f\}$. The optimal fertility choice ($b^{*}$) is then given in a closed-form function. One can confirm that the sign of the derivative of optimal fertility with respect to the wage of female ($\frac{\partial b^{*}}{\partial w_{f}}$) is negative under the following condition:
@@
Online Appendix B provides details of the derivation. Under a paternalistic social norm stipulated in the Old Civil Code, the head’s housework productivity, particularly childrearing, was negligible ($\kappa_{m}=0$). It follows that condition 1 implies $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$ for my empirical setting. Since the marketization of childcare ($\gamma$) was considerably low at that time, the optimal fertility was also negatively associated with the relative wage ($w_{f}/w_{m}$). [^19] As mechanization progressed, miners were required to accumulate knowledge and experience in machinery operation. Management, therefore, had an incentive to encourage miners to form families and continue employment over a long term (Ogino 1993). Thus, many coal mines were equipped with nurseries and other welfare facilities from the 1920s (Ministry of Commerce and Industry 1926). However, evidence shows that these nurseries required fees and had covered only a small part of their childcare requirements. See Online Appendix B for details. This result suggests the following two predictions:
@@
  1. Prediction 1: Before female labor regulation, the mining area could have lower fertility, given the higher relative wages in the mining sector.
@@
  2. Prediction 2: After female labor regulation, the mining area could have recovered fertility in response to the reduction in the relative wages.
@@
The short-run population dynamics in a local economy depend on natural and social changes. [^20] From year $t$ to $t+1$, this geometric population growth model is specified as $\Psi_{t+1}-\Psi_{t}=(\Upsilon_{t}-\Theta_{t})+\Pi_{t}$, where $\Psi$, $\Upsilon$, $\Theta$, and $\Pi$ indicate the number of people, live births, deaths, and net migrants, respectively. In the former change, the lower fertility under the substitution effect before the regulation would not be a major factor in population growth in the mining area. In turn, the fertility gain after the regulation could increase population growth. Mortality term in the natural change is given outside of the household’s fertility choice in the short-run. The regulation of female miners reduced their occupational hazards, which could improve female mortality. The improvements in maternal health could also mitigate prenatal and neonatal deaths (Section 2.3). Both ameliorations would work to reduce population losses in the mining area. The influence of social change on local population growth is uniquely determined when the sign of natural change is determined. Moreover, social change is positive in the mining area because the regional wage gap would encourage the influx of workers from surrounding agrarian areas into the mining area. For female workers, the wage reduction via the regulation could reduce the attractiveness of coal mining labor. However, it may not offset the benefits for couples moving into the mining area because male miners’ wages were not reduced by the regulation. [^21] Thus, since female workers migrated with their husbands, the migration of couples is less affected by the regulation than that of single workers. To summarize, while the mining area had an influx of workers throughout the period under consideration, the fertility and mortality in the mining area could be affected by the enforcement of the revised regulation of 1933.
@@ 4 Data
This study created a unique dataset of the demographics across mining areas in the Japanese archipelago. As mining is a localized economic activity, a smaller lattice dataset is preferred to identify the impact of mines. A set of official census-based municipal-level statistics published in the 1930s was collected and digitized.
@@ 4.1 Mine Deposit
I use official reports named Zenkoku kōjyō kōzan meibo (lists of factories and mines) published by Association for Harmonious Cooperation (1932, 1937) (hereafter, called the AHC), which document coal mine locations measured in October 1931 and October 1936. The AHC listed all mines employing $50$ miners and more. Thus, the main target of this study is the average impact of small- to large-scale coal mines and not the micro-scale collieries employing fewer than $50$ mimers. [^22] Generally, several coal mines are clustered in a coal mining municipality. For example, 85 (108) out of 93 (115) municipalities with mines had one to four mines in the 1930 (1935) sample. In this sense, this study estimates the average impact of these general coal mining municipalities. In Online Appendix D.11, I confirmed that the main findings are not sensitive to a small number of municipalities with several mines. Despite this, the AHC is still comprehensive as it includes information on small-scale mines, which have been neglected in the literature.
The treatment group is defined as the municipalities located within 0–5 km of the centroid of a municipality with mines. [^23] The study used an official shapefile provided by the Ministry of Land, Infrastructure, Transport and Tourism for geocoding. See Online Appendix C.1 for the details. A 5 km threshold is fundamentally plausible because the mean value of the distance to the nearest neighborhood municipality is approximately 4 km. In addition, the estimated effects disappeared in the statistical sense outside the 5 km range, which means that the potential spillover effect is captured under this threshold. Online Appendix D.6 summarizes finer details of the validity of the threshold. For the control group, the threshold is set as 5–30 km from the centroid. [^24] The definition of both groups is conservative compared with that in the literature, which uses 10–20 km and 100–200 km as the thresholds of the treatment and control groups, respectively (Wilson 2012; Aragón and Rud 2015; Kotsadam and Tolonen 2016; Benshaul-Tolonen 2019). This is because Japan has a smaller and thinner archipelago than countries studied in the literature, such as South Africa (Online Appendix Figure 2). In Online Appendix D.6, further evidence is provided that 5–30 km is a plausible distance for the control group given that the impacts of the coal mines are concentrated within 5 km distance from the mines.
@@
Figure 3: Location of Coal Mines around Chikuhō Coalfield in Kyūshū
Notes: The white circles indicate the location of municipalities with coal mines around Chikuhō coalfield (Figure 2a) in Kyūshū region. Figure 3a shows the coal mines in Chikuhō, Kasuya, and Miike coalfields in Fukuoka prefecture, Karatsu coalfield in Saga prefecture, and Miike coalfield in Kumamoto prefecture. Amakusa coalfield in Kumamoto prefecture is added in Figure 3b (on the lower left of the figure). Ōita prefecture does not have coal mines but is shown in the figures to explain the border of the sample. Treatment (control) group includes municipalities within 5 (between 5 and 30) km from a mine. The excluded municipalities are shown as empty lattices in the figures.
@@
Figure 2 indicates the coal mine locations from the AHC and the corresponding treatment and control groups by year. Most coal mines are located in three specific areas in the southernmost, central, and northeastern regions. These are representative coalfields called Chikuhō, Jyōban, and Ishikari-Kushiro Coalfields, respectively. Notably, a comparison of Figures 2a and 2b shows that these agglomerations were stable over time. For a closer look, Figure 3 provides an example map for the most representative coalfield called Chikuhō Coalfield in Kyūshū, Japan’s southernmost island (indicated in Figure 2a). One can observe that the distribution of mines has changed little, with only a few mines either opened or closed between 1931 and 1936. This feature of coal mines makes it difficult to use within-variation in the spatiotemporal distribution of the mines for identification. Online Appendix D.8 discusses this point in detail. [^25] In short, it is not feasible to implement a difference-in-differences estimator for the full sample because of the lack of within-variation (i.e., the number of municipalities that experienced the opening/closing of the coal mines). Nevertheless, in Online Appendix D.8, the results from the two-group by two-period difference-in-differences setting for a small subsample including Ogawa and Miyoshiyama coal mines will be demonstrate (Figure 2b), which provide the results reasonable to the findings.
Finally, the generation of the analytical samples will be explained. From $11,151$ total municipalities, the municipalities within a $30$ km radius from the centroid of the municipalities with mines are first retained. The intersection of the different types of mines are then excluded, such as municipalities within $5$ km from a gold, silver, and copper (GSC) mine. Online Appendix Table C.3 presents the summary statistics of the treatment variables. The number of municipalities in the analytical sample based on the 1931 and 1936 AHC were $1,140$ and $1,364$, respectively. Meanwhile, the proportion of treated municipalities was approximately $15$% and $13$%, respectively, for each sample. [^26] The number of municipalities with coal mines in each analytical sample is 93 and 115, respectively. The slight reduction in the share of the treatment group is due to the increase in the number of enclaves, which is explained above. A mine opening in the enclave area increases the number of controlled (i.e., surrounding) municipalities, which leads to a decrease in the relative share of the treatment group. Examples are the Ogawa and Miyoshiyama coal mines in Tōhoku region (Figure 2b).
@@ 4.2 Geological Stratum
The spatial distribution of a specific geological stratum created in the Cenozoic era, which includes some carboniferous ages, is used as an instrumental variable (IV) for the location of coal mines (Section 5.1). Data on the geological strata are obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism, which includes roughly nine thousand stratum points. Each municipality in the analytical sample was matched with the nearest stratum point to identify the municipalities with the relevant stratum. Online Appendix C.2 provides finer details of the definition and an example of a geological columnar section of a representative coal mine to show the relevance of the IV. Online Appendix Table C.3 presents the summary statistics of the indicator variables for the stratum.